# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v2

@@ 4.5 Linear vs Nonlinear Models with Causal Predictors: Case Study
To illustrate the role of causally relevant predictors, we present Hurricane Larry (2021) as a case study. We selected Larry because, among the storms in the independent test set, it showed the largest increase in predictive skill (R²) when using SHIPS+ MLPs with causal predictors compared to the operational SHIPS model. This allows us to investigate which features and causal predictors contributed most to the improvement in forecast performance. Larry was a long-lived and intense Cape Verde hurricane that underwent a period of RI over the tropical Atlantic before eventually making landfall in Newfoundland as a Category 1 hurricane. For this analysis, we truncate the time series before its extratropical transition, focusing solely on the tropical phase. Figure 7a shows the best track of Larry during the evaluation period (Left), and a comparison of model performance (Right) between MLR and MLP, both with and without the addition of causally selected predictors. MLR SHIPS and MLR SHIPS+ both generally under-predict intensity change throughout the period, showing less variability and weaker response to fluctuations. Both versions of the Multi-Layer Perceptron (MLP) models with SHIPS and SHIPS+ (causally selected features) are consistently closer to the ground truth (black line) than the corresponding MLR models.
@@
Figure 7: Hurricane Larry (2021): (a) Best track from IBTrACS showing the time used for testing (b) performance comparison of MLR vs. MLP comparing IBTrACS 24 hour intensity change (truth) to predictions with (SHIPS+) and without causal predictors (SHIPS). Overall, the MLP consistently outperforms the MLR, with the addition of predictors improving the performance notable from 54 hours. (c) Lifetime-maximum absolute SHAP values highlighting the dominant role of the added predictors (R001, PVOR). (d) Time-series decomposition of $\Delta y_{\mathrm{pred}}$ showing R001 and PVOR as the main contributors, with smaller adjustments from LAT and POT.
@@
MLP SHIPS generally over-predicts TC Larry’s intensification, except between Hour 24-48. Adding SHIPS+ dampens this overestimation, bringing it closer to the true signal, especially after 54 hours. In contrast, the performance gain from adding causal variables to MLR is minimal, indicating that linear models cannot fully leverage the additional causal features, highlighting the limitations of linear regression for dynamic systems like TCs, while the MLP models show a more dynamic response to intensity fluctuations.
@@
To assess the role of causal predictors, we test whether the SHIPS++ MLP’s gains arise from learning new relations for the added predictors or simply from reweighting existing SHIPS predictors. Therefore, we apply the SHAP-based decomposition of Grundner et al. (2022) (Appendix B) to compare the SHIPS model $f$ and the SHIPS++ model $g$. By SHAP additivity, the prediction difference can be decomposed as follows:
[…]
@@
For Hurricane Larry, lifetime-mean absolute SHAP values (Fig.7c) indicate that Eq.3 is well approximated by the contributions of two added predictors (R001 and PVOR) plus adjustments to two existing predictors (LAT and POT). The time-series decomposition (Fig.7d) confirms that R001 and PVOR are the two most important contributors to $\Delta y_{\mathrm{pred}}$, with smaller adjustments from LAT and POT and a small residual from all other features over the period of interest (30-120 hour). This lends credence to the interpretation that SHIPS++ improvements primarily arise from the added causal predictors. Consistent with the partial dependence for R001 (Fig.6e), SHIPS underestimates the weakening of Larry after Hour 54 because the original predictors under-represent the suppressing influence of inner-domain near-surface humidity; including R001 mitigates this bias.
@@ 4.6 Testing the new predictors in operational SHIPS
As in SHIPS development, candidate variables are added to the developmental set and undergo a predictor-screening test for statistical significance with respect to intensity changes at 6 hour increments to lead times of 168 h. Training follows the SHIPS perfect-prognosis protocol: predictors are averaged over each forecast interval and sampled from GFS analysis fields about the future best-track positions at the valid times. In operations, the same predictors are computed from GFS forecast fields along the National Hurricane Center (NHC) forecast track. Note that the causal predictor identification used in this manuscript (Table 1) differs by using only $t=0$ values about the current best-track position—by design, to keep feature discovery in a purely statistical forecasting (not post-processing) setting and to make the causal tests consistent with the regression task. We then apply the standard SHIPS screening and coefficient-update procedure described next. Three conditions are needed to pass the screening step: (1) When added one at a time, a new predictor must increase the variance explained by the model by at least 0.2% averaged over 5 consecutive forecast intervals; (2) The regression coefficient must be significant at the 99% level for at least 5 forecast intervals; (3) All predictors that pass steps 1 and 2 are added to the training sample and the coefficients are recalculated. The regression coefficients for each predictor must still be significant at the 99% level for at least five forecast intervals. If a predictor does not pass step 3, the least significant predictor is removed and then step 3 is repeated until all retained predictors pass the significance test. Results showed that SHMD, R001 and PVOR passed steps 1 and 2 for the 1982-2021 North Atlantic training sample. When all three were added, they also passed step 3. The next step in testing new SHIPS predictors is to perform retrospective forecasts on independent cases using only the input that is available in real time (forecast tracks and GFS forecast fields). To allow for a fair comparison, SHIPS was trained on the 1982-2021 sample with the original 28 operational predictors (baseline) and then with the addition of the three new predictors and both versions were run on the 2022-2024 independent cases. This sample included 794 cases with a 12 hour forecast, which decreased to 114 cases by 168 hours since many TCs dissipate by seven days. The mean absolute error (MAE) of the intensity forecasts for the baseline SHIPS and with the three new predictors were calculated using the intensity in the final NHC best track as “truth” following the standard NHC forecast procedure Cangialosi et al. (2024). The verification sample includes tropical and subtropical cyclones but excludes the extratropical and pre-genesis stages.
@@
Figure 8 shows the percent improvement (reduction in MAE) of SHIPS with the new predictors relative to baseline. SHIPS forecasts improved at all forecast times, with the largest improvements at longer forecast times. The relatively small improvements at short lead times reflect the dominant role of the persistence term in operational SHIPS, together with strong constraints imposed by initial conditions and directly assimilated observations, which limit the contribution of GFS-based environmental predictors. As forecast lead time increases, the influence of persistence and initial conditions weakens, while large-scale environmental controls become more important, allowing the newly identified predictors to produce larger gains in forecast skill. The improvements at 120-168 hours were statistically significant at the 90% level using a standard statistical test that accounts for serial correlation. Although causal discovery used only $t=0$ values, the longer-lead time gains seen in Figure 8 likely reflect the statistical–dynamical use of forecast fields in operations, motivating extension of the discovery procedure to predictors evaluated during the forecast period. The three new predictors (SHMD, R001 and PVOR) will be considered for implementation in future operational versions of SHIPS.
@@ 5 Conclusions
To improve operational SHIPS, we introduced a multidata PC causal discovery framework to discover causally relevant predictors that drive North Atlantic TC intensity changes. First, we conducted tests by replicating the SHIPS predictors with the ERA5 reanalysis: we expanded the predictor list by testing combinations of dynamic and thermodynamic variables across different vertical layers and storm radial areas. M-PC tests using this augmented dataset revealed additional predictors that consistently demonstrated clear links to TC intensity change. Compared to other feature selection baselines, the causal approach performed best for short-range lead times (24–72 hours), highlighting its ability to isolate physically meaningful predictors while filtering out spurious relationships. Through this process, we shortlisted six key variables that were not currently used as operational SHIPS predictors. One of the new predictors, the surface level relative humidity for the outer area of the storm (R000), stands out because it is (1) already part of the extended SHIPS predictors list; and (2) has been linked to tropical cyclone intensification and RI across multiple basins Wu et al. (2012); Richardson (2020); Wang et al. (2020). The inclusion of other predictors, such as low-level shear and mid-tropospheric potential vorticity, underscores the importance of multiscale dynamical interactions and moisture–vorticity coupling processes that are often missed by traditional statistical frameworks.
@@
The Hurricane Larry (2021) case study illustrated the added value of causally relevant predictors. Although both MLR and MLP models benefit from causal features, the nonlinear MLP better leverages these predictors to capture dynamic intensity fluctuations, including RI and decay phases, even though the predictors were selected using a linear causal discovery procedure. In contrast, linear models tend to under-predict variability and respond less effectively to short-term changes, highlighting the advantage of nonlinear approaches in capturing complex TC behavior. These findings are consistent with the operational SHIPS tests, showing improvements at all lead times and the largest MAE reductions at 120–168 hour (Fig.8).
While purely statistical models already benefit from causal feature selection and nonlinear methods, our results also point to their limitations at longer lead times, where static predictors become less informative. In addition, the causal discovery and operational validation are constrained by the resolution and variable availability of ERA5 and GFS analyses, which may limit the representation of inner-core processes and exclude physically relevant predictors such as surface latent heat fluxes. However, in a statistical-dynamical framework such as operational SHIPS, where the GFS forecast model provides the evolving environmental fields, these limitations can diminish, as evidenced by the largest improvements at 120–168 h. A promising next step is to derive causally selected predictors directly from dynamical forecast output and along forecast trajectories. This would allow their seamless integration into statistical-dynamical frameworks such as SHIPS, potentially improving the forecast skill for RI and addressing one of the main sources of uncertainty in operational TC intensity prediction.
@@ Acknowledgments
We thank Dr. Andreas Gerhardus (DLR, Germany) for the initial guidance on causal experiments with replicated data. The authors also acknowledge the DCSR at UNIL for providing computational resources and technical support. We thank the two anonymous reviewers for their detailed reviews and constructive suggestions. Tom Beucler acknowledges support from the Swiss National Science Foundation (SNSF) under Grant No. 10001754 (“RobustSR” project).
@@
We confirm that this study is original research, developed entirely by the authors, and has been submitted to Weather and Forecasting.

### v3

@@ 4.5 Linear vs Nonlinear Models with Causal Predictors: Case Study
To illustrate the role of causally relevant predictors, we present Hurricane Larry (2021) as a case study. We selected Larry because, among the TCs in the independent test set, it showed the largest increase in predictive skill (R²) when using SHIPS+ MLPs with causal predictors compared to the operational SHIPS model. This allows us to investigate which predictors drive the improvement in forecast performance. Larry was a long-lived and intense Cape Verde hurricane that underwent a period of RI over the tropical Atlantic before eventually making landfall in Newfoundland as a Category 1 hurricane. For this analysis, we truncate the time series before its extratropical transition, focusing solely on the tropical phase. Fig.  7 shows the best track of Larry during the evaluation period (a), and a comparison of model performance (b) between MLR and MLP, both with and without the addition of causally selected predictors. MLR SHIPS and MLR SHIPS+ both generally under-predict intensity change throughout the period, showing less variability and weaker response to fluctuations. Both versions of the Multi-Layer Perceptron (MLP) models with SHIPS and SHIPS+ (causally selected features) are consistently closer to the ground truth (black line) than the corresponding MLR models.
@@
Figure 7: Hurricane Larry (2021): (a) Best track from IBTrACS showing the time used for testing (b) performance comparison of MLR vs. MLP comparing IBTrACS 24-hour intensity change (truth) to predictions with (SHIPS+) and without causal predictors (SHIPS). Overall, the MLP consistently outperforms the MLR, with the addition of predictors improving the performance notable from 54 hours. (c) Lifetime-maximum absolute SHAP values highlighting the dominant role of the added predictors (R001, PVOR). (d) Time-series decomposition of $\Delta y_{\mathrm{pred}}$ showing R001 and PVOR as the main contributors, with smaller adjustments from LAT and POT.
@@
MLP SHIPS generally over-predicts TC Larry’s intensification, except between 24 to 48 hours. Adding SHIPS+ dampens this overestimation, bringing it closer to the true signal, especially after 54 hours. In contrast, the performance gain from adding causal variables to MLR is minimal, indicating that linear models cannot fully leverage the additional causal features, highlighting the limitations of linear regression to capture TC dynamics, while the MLP models show a more dynamic response to intensity fluctuations.
@@
To assess the role of causal predictors, we test whether the SHIPS++ MLP’s gains arise from learning new relations for the added predictors or simply from reweighting existing SHIPS predictors. Therefore, we apply the SHAP-based decomposition of (Grundner et al., 2022) (Appendix B) to compare the SHIPS model $f$ and the SHIPS++ model $g$. By SHAP additivity, the prediction difference can be decomposed as follows:
$\begin{split}\Delta y_{\mathrm{pred}}=B_{f}-B_{g}&+\underbrace{\sum_{i=1}^{N_{\mathrm{SHIPS}}}\bigl(\mathrm{SHAP}_{i}^{f}-\mathrm{SHAP}_{i}^{g}\bigr)}_{\text{common predictors}}\\ &-\underbrace{\sum_{i=1}^{N_{\mathrm{causal}}}\mathrm{SHAP}_{i}^{g}}_{\text{added (causal) predictors}}\,\end{split}$ (3)
@@
For Hurricane Larry, lifetime-mean absolute SHAP values (Fig. 7c) indicate that Eq.3 is well approximated by the contributions of two added predictors (R001 and PVOR) plus adjustments to two existing predictors (LAT and POT). The time-series decomposition (Fig. 7d) confirms that R001 and PVOR are the two most important contributors to $\Delta y_{\mathrm{pred}}$, with smaller adjustments from LAT and POT and a small residual from all other features over the period of interest (30-120 hours). This lends credence to the interpretation that SHIPS++ improvements primarily arise from the added causal predictors. Consistent with the partial dependence for R001 (Fig. 6e), SHIPS underestimates the weakening of Larry after 54 hours because the original predictors under-represent the suppressing influence of inner-domain near-surface humidity; including R001 mitigates this bias.
@@ 4.6 Testing the new predictors in operational SHIPS
As in SHIPS development, candidate variables are added to the developmental set and undergo a predictor-screening test for statistical significance with respect to intensity changes at 6-hour increments to lead times of 168 hours. Training follows the SHIPS perfect-prognosis protocol: predictors are averaged over each forecast interval and sampled from GFS analysis fields about the future best-track positions at the valid times. In operations, the same predictors are computed from GFS forecast fields along the National Hurricane Center (NHC) forecast track. Note that the causal predictor identification used in this manuscript (Table 1) differs by using only $t=0$ values about the current best-track position by design, to keep feature discovery in a purely statistical forecasting (not post-processing) setting and to make trse causal tests consistent with the regression task. We then apply the standard SHIPS screening and coefficient-update procedure described next. Three conditions are needed to pass the screening step: (1) When added one at a time, a new predictor must increase the variance explained by the model by at least 0.2% averaged over 5 consecutive forecast intervals; (2) The regression coefficient must be significant at the 99% level for at least 5 forecast intervals; (3) All predictors that pass steps 1 and 2 are added to the training sample and the coefficients are recalculated. The regression coefficients for each predictor must still be significant at the 99% level for at least five forecast intervals. If a predictor does not pass step 3, the least significant predictor is removed and then step 3 is repeated until all retained predictors pass the significance test. Results showed that SHMD, R001 and PVOR passed steps 1 and 2 for the 1982-2021 North Atlantic training sample. When all three were added, they also passed step 3. The next step in testing new SHIPS predictors is to perform retrospective forecasts on independent cases using only the input that is available in real time (forecast tracks and GFS forecast fields). To allow for a fair comparison, SHIPS was trained on the 1982-2021 sample with the original 28 operational predictors (baseline) and then with the addition of the three new predictors and both versions were run on the 2022-2024 independent cases. This sample included 794 cases with a 12-hour forecast, which decreased to 114 cases by 168 hours since many TCs dissipate by seven days. The mean absolute error (MAE) of the intensity forecasts for the baseline SHIPS and with the three new predictors were calculated using the intensity in the final NHC best track as “truth” following the standard NHC forecast procedure Cangialosi et al. (2024). The verification sample includes tropical and subtropical cyclones but excludes the extratropical and pre-genesis stages.
@@
Fig.  8 shows the percent improvement (reduction in MAE) of SHIPS with the new predictors relative to baseline. SHIPS forecasts improved at all forecast times, with the largest improvements at longer forecast times. The relatively small improvements at short lead times reflect the dominant role of the persistence term in operational SHIPS, together with strong constraints imposed by initial conditions and directly assimilated observations, which limit the contribution of GFS-based environmental predictors. As forecast lead time increases, the influence of persistence and initial conditions weakens, while large-scale environmental controls become more important, allowing the newly identified predictors to produce larger gains in forecast skill. The improvements at 120-168 hours were statistically significant at the 90% level using a standard statistical test that accounts for serial correlation. Although causal discovery used only $t=0$ values, the longer-lead time gains seen in Fig.  8 likely reflect the statistical–dynamical use of forecast fields in operations, motivating extension of the discovery procedure to predictors evaluated during the forecast period. The three new predictors (SHMD, R001 and PVOR) will be considered for implementation in future operational versions of SHIPS.
@@ 5 Conclusions
To improve operational SHIPS, we introduced a multidata PC causal discovery framework to discover causally relevant predictors that drive North Atlantic TC intensity changes. First, we conducted tests by replicating the SHIPS predictors with the ERA5 reanalysis: we expanded the predictor list by testing combinations of dynamic and thermodynamic variables across different vertical layers and storm radial areas. M-PC tests using this augmented dataset revealed additional predictors that consistently demonstrated clear links to TC intensity change. Compared to other feature selection baselines, the causal approach performed best for short-range lead times (24–72 hours), highlighting its ability to isolate physically meaningful predictors while filtering out spurious relationships. Through this process, we shortlisted six key variables that were not currently used as operational SHIPS predictors. One of the new predictors, the surface level relative humidity for the outer area of the TC (R000), stands out because it is (1) already part of the extended SHIPS predictors list; and (2) has been linked to TC intensification and RI across multiple basins Wu et al. (2012); Richardson (2020); Wang et al. (2020). The inclusion of other predictors, such as low-level shear and mid-tropospheric potential vorticity, underscores the importance of multiscale dynamical interactions and moisture–vorticity coupling processes that are often missed by traditional statistical frameworks.
@@
The Hurricane Larry (2021) case study illustrated the added value of causally relevant predictors. Although both MLR and MLP models benefit from causal features, the nonlinear MLP better leverages these predictors to capture dynamic intensity fluctuations, including RI and decay phases, even though the predictors were selected using a linear causal discovery procedure. In contrast, linear models tend to under-predict variability and respond less effectively to short-term changes, highlighting the advantage of nonlinear approaches in capturing complex TC behavior. These findings are consistent with the operational SHIPS tests, showing improvements at all lead times and the largest MAE reductions at 120–168 hours (Fig. 8).
Although purely statistical models already benefit from causal feature selection and nonlinear methods, our results also point to their limitations at longer lead times, where static predictors become less informative. In addition, causal discovery and operational validation are constrained by the resolution and variable availability of ERA5 and GFS analyses, which may limit the representation of inner-core processes and exclude physically relevant predictors such as surface latent heat fluxes. However, in a statistical-dynamical framework such as operational SHIPS, where the GFS forecast model provides the evolving environmental fields, these limitations can diminish, as evidenced by the largest improvements at 120–168 hours. A promising next step is to derive causally selected predictors directly from dynamical forecast output and along forecast trajectories. This would allow their seamless integration into statistical-dynamical frameworks such as SHIPS, potentially improving the forecast skill for RI and addressing one of the main sources of uncertainty in operational TC intensity prediction.
@@ Acknowledgments
The authors thank Dr. Andreas Gerhardus (DLR, Germany) for the initial guidance on causal experiments with replicated data. The authors also acknowledge the DCSR at UNIL for providing computational resources and technical support. We thank the two anonymous reviewers for their detailed reviews and constructive suggestions. Tom Beucler acknowledges support from the Swiss National Science Foundation (SNSF) under Grant No. 10001754 (“RobustSR” project).
@@
We confirm that this study is original research, developed entirely by the authors, and has been submitted to Weather and Forecasting (AMS).